P 3-30 (similar to Question Help Your grandfather put some money into an account for you on the day you were born. You are now 18 years old and are allowed to withdraw the money. The account currently has $8,512 in it and pays an interest rate of 10% a. How much money would be in the account if you left the money there until your 25th birthday? b. What if you left the money until your 65th birthday? c. How much money did your grandfather originally put into the account? a. How much money would be in the account if you left the money there until your 25th birthday? The amount that would be in the account if you left the money there until your 25th birthday would be (Round to the nearest dollar.)You are thinking of building a new machine that will save you $3,000 in the first year. The machine will then begin to wear out so that the savings decline at a rate of 1% per year forever. What is the present value of the savings if the interest rate is 7% per year? The present value of the savings is $나 (Round to the nearest dollar.)
Answer #1

1. a) Number of years = 7
Rate = 10%
FV = 8512*(1+10%)7 = 16587.48

b) Money left after 65th year = 8512*(1+10%)47 = 750,736.99

c) The amount your grandfather put into money = 8512/(1+10%)18 = 1530.96

Max 1 question can be solved.

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well